OpenStudy (anonymous):

Write down the following in order of size,smallest first. cos100 sin100 tan100

OpenStudy (sidsiddhartha):

use a calculator

OpenStudy (vishweshshrimali5):

I hope that 100 here is in radians

OpenStudy (vishweshshrimali5):

cos(100) would be the largest as 100 radians lies in 4th quadrant and the only +ve function among all.

OpenStudy (vishweshshrimali5):

|sin(100)| < |tan(100)| but since both are negative. So, sin(100) > tan(100)

OpenStudy (vishweshshrimali5):

Thus, my preliminary answer would be: cos(100)>sin(100)>tan(100) And great I am correct

OpenStudy (vishweshshrimali5):

I am actually surprised :D

OpenStudy (anonymous):

I believe 100 is in degree

OpenStudy (anonymous):

cos 100 degree is less than 0

OpenStudy (anonymous):

sin 100 is just under 1

OpenStudy (anonymous):

tan 100 is almost - infinity

OpenStudy (anonymous):

so tan 100<<cos100<sin100

OpenStudy (vishweshshrimali5):

But usually in a question if degree symbol is not used then it is assumed that the angle is in radians. After all, SI Unit is radians not degrees.

OpenStudy (anonymous):

but for this case if it is in degree the sizes can be guessed. If it is in radians then we need to consult the calculator. I believe the question calls for understanding of the behaviour of the cos, sin and tan curves.

OpenStudy (vishweshshrimali5):

well even if it is in radians you can solve it without using calculator... except for division and multplication

OpenStudy (anonymous):

ok thanks

OpenStudy (haseeb96):

@sheelo_mughal listen i am going to tell one thing which surprise u

OpenStudy (haseeb96):

listen sine function is increase from 0 to 90 and cos function decrease 0 to 90 and tan function also increase from 0 to 90 u can do it on yur calculator or from the trigonometric function like sin 30 = 0.5 sin 45 = 0.70 sin 60 = 0.866 sin 90 = 1 here it is showing that function is increasing so now u can u can found that first cos 100 will come first then sin and after it tan will come

OpenStudy (vishweshshrimali5):

@Haseeb96 , are you assuming the angles to be radian or in degrees ?

OpenStudy (haseeb96):

yes

OpenStudy (vishweshshrimali5):

radians ?

OpenStudy (haseeb96):

no degrees

OpenStudy (vishweshshrimali5):

Then sorry but cos(100) would not be maximum

OpenStudy (vishweshshrimali5):

As it will be negative and sin(100) would be positive

OpenStudy (vishweshshrimali5):

So, sin(100 degrees) > cos(100 degrees)

OpenStudy (vishweshshrimali5):

Correct order would be: tan(100 deg) < cos(100 deg) < sin(100 deg)

OpenStudy (vishweshshrimali5):

The overall answer is: \[\large{\boxed{\tan(100^\circ) < \cos(100^\circ) < \sin(100^\circ)}}\] and \[\large{\boxed{\cos(100)>\sin(100)>\tan(100)}}\]

OpenStudy (vishweshshrimali5):

Verified using calculator ^^^